OpenStudy (brittany16):

Which operation(s) would isolate the variable in this equation? 3y - 12 = 14

OpenStudy (anonymous):

ADD 12 DIVIDE BY 3 Y=3

OpenStudy (anonymous):

Y=12 SORRY

OpenStudy (anonymous):

add 12 to both sides 3y=26 divide both sides by 3=26/3

OpenStudy (anonymous):

add 12 and then divide by 3 3y-12=14 3y=26 y=8 and 2/3

jimthompson5910 (jim_thompson5910):

3y - 12 = 14 3y - 12 + 12 = 14 + 12 ... Add 12 to both sides 3y = 26 3y/3 = 26/3 ... Divide both sides by 3 y = 26/3